# Auto-Calibration and Biconvex Compressive Sensing with Applications to Parallel MRI

### v1

@@ 4.1 Assumptions and main theorem
Before we move on to present our main theoretical findings, for convenience we summarize all the assumptions on the model (14) and the signals here:
  1. 1.
    The rows of B are chosen uniformly at random without replacement from the rows of a fixed matrix $\textbf{B}_{0}$ and $\textbf{B}_{0}^{*}\textbf{B}_{0}=\textbf{I}_{k}$.
  2. 2.
    $\mathbf{\Psi}$ is a known and fixed orthonormal basis with $\mathbf{\Psi}^{*}\mathbf{\Psi}=\mathbf{\Psi}\mathbf{\Psi}^{*}=\textbf{I}$.
  3. 3.
    The rows of $\textbf{F}_{\Omega}$ are chosen uniformly at random without replacement from the DFT matrix with $\textbf{F}^{*}\textbf{F}=\frac{N}{L}\textbf{I}_{N}$.
  4. 4.
    We analyze two probability models for the sensitivity parameters:
@@
Theorem 1
[…]
$\|\textbf{X}-\textbf{X}_{0}\|_{F}\leq C_{1}\|\mathcal{P}_{S^{c}}\textbf{X}_{0}\|_{1,2}+(C_{2}+C_{3}\sqrt{s})\sigma$ (21)
with probability $1-4N^{-\alpha}$ if
  1. 1.
    the number of measurement satisfies
    $L\geq C_{\alpha}kn{\mu^{B}_{max}}^{2}{\mu^{\Psi}_{max}}^{2}\log(knN).$
  2. 2.
    If $\textbf{h}_{l}$ is independently chosen from a basis or lie on the complex sphere and the number of coils satisfy
    $C\geq C_{\alpha}\log(N)k.$
Here, $\mu^{B}_{max}:=\max\limits_{i,j}\sqrt{L}|\mathbf{B}_{i,j}|$, $\mu^{\Psi}_{max}:=\max\limits_{i,j}\sqrt{N}|{\boldsymbol{\Psi}}_{i,j}|$, which measures the incoherence of the rows of B and $\boldsymbol{\Psi}$ respectively. The constant $C_{\alpha}$ grows linearly with respect to $\alpha$. A detailed bound for the constants can be found in the corresponding proofs, see Section 5.
Remark: The recovery is stable and robust. Note that in the bound, $kn={{\|P_{S}(\textbf{X}_{0})\|}_{0}}$ and when $S$ is the true support of $\textbf{z}_{0}$, $\mathcal{P}_{S^{c}}\textbf{X}_{0}=0$. For the bound regarding $L$, we expect that the magnitude of each entry in $\mathbf{B}$ and $\boldsymbol{\Psi}$ does not vary too much for it to be useful. For example, when the columns of $\mathbf{B}$ and $\boldsymbol{\Psi}$ are chosen from DFT matrices, we optimally have $\mu^{B}_{max}=1$, $\mu^{\Psi}_{max}=1$.
@@
We would like to develop a non-uniform recovery result to take advantage of the randomness introduced by using more parallel measurements. We first comment on our choice of assumptions. The first and third assumptions are on the sensing matrix. Note that B is chosen uniformly at random from all possible row permutations of a fixed matrix $\textbf{B}_{0}$. The orthogonality of the columns of B is not affected by row permutations since $\textbf{B}_{i}^{*}\textbf{B}_{j}=\sigma(\textbf{B}_{i})^{*}\sigma(\textbf{B}_{j})$. Consider when the rows of B are randomly permuted. For any h fixed, we have the coil sensitivity $\mathbf{s}=\textbf{B}\textbf{h}$ to be randomly permuted copies of some underlying $\textbf{s}_{0}=\textbf{B}_{0}\textbf{h}$. On the other hand, if we assume no randomness in B, if there are zero entries in $\textbf{s}=\textbf{B}\textbf{h}$, it is impossible to recover uniquely the values of the image z at those locations. The randomness helps in providing an average case analysis for each coil. The third assumption is a standard one in compressive sensing theory on the random sampling scheme for k-space/Fourier space measurements. The assumption on H is critical for developing an average cage analysis to explain the improved results when more parallel measurements are taken.
Ideally, we would like to derive a proof for the setup when both B and $\mathbf{\Psi}$ are fixed, and the randomness comes from the support set $S$ of z alone. The result would be analogous to the conditioning of the random sub-matrices of a fixed sampling matrix, e.g. see Chapter 14 of [21]. Yet, we encountered some unsolvable difficulties in finding a realistic bound for the incoherence measure between the sub-blocks of the sampling matrix A. Nevertheless, we still provide a formal statement and some discussions in the Appendix 7. The main theorem enables self-calibration and ensures stable and robust recovery in the presence of sparsity defects and noise.
@@
Next, let us discuss the reason of including multiple measurements/coils and the probability model on H. Consider the worst case scenario where all the $\textbf{h}_{i}$ are identical, then the measurements $\textbf{Y}=\textbf{A}\textbf{X}$ consist of identical columns. By including multiple measurements, we are not gaining more information on X regardless of the number of repeated measurements taken. Hence, employing multiple measurements does not necessarily outperform the single measurement case. The following proposition gives a formal statement. The proof follows exactly from Proposition 4.1 in [19].
@@
However, as observed in many real world applications, one would expect multiple measurements to provide more information that should aid the recovery tasks in mind. In an MRI setting, it has been shown that parallel imaging (pMRI) outperforms single-coil MRI in many cases, and the incoherent coils are often used as an explanation for the improved results. To understand the role of including multiple measurements, and from the worst case analysis, it suggests a need to impose some probability model on the signals to be recovered in order to prove some average case results.
We consider the underlying image $\textbf{z}=\mathbf{\Psi}\textbf{x}$ to be fixed and only impose some probability distribution on $\textbf{h}_{i}$. That is to say we are imaging a fixed object with varying coil sensitivities where the spanning coefficient of the coils follows a certain probability distribution. For the two choices on the probability models on H, they all ensure incoherence between the coils and that the coil sensitivities have the same energy. Observe first that, the incoherence between coils $\textbf{s}_{i}^{*}\textbf{s}_{j}=(\textbf{B}\textbf{h}_{i})^{*}(\textbf{B}\textbf{h}_{j})=\textbf{h}_{i}^{*}\textbf{B}^{*}\textbf{B}\textbf{h}_{j}=\textbf{h}_{i}^{*}\textbf{h}_{j}$ is small for $k$ large, and the coils all have the same energy $\textbf{s}_{i}^{*}\textbf{s}_{i}=(\textbf{B}\textbf{h}_{i})^{*}(\textbf{B}\textbf{h}_{i})=\textbf{h}_{i}^{*}\textbf{B}^{*}\textbf{B}\textbf{h}_{i}=\textbf{h}_{i}^{*}\textbf{h}_{i}=1$.
@@
The proof of the main theorem follows a well-established proof outline of compressive sensing [21]. First, we refer to a sufficient condition from [20] to ensure a robust and stable recovery. Subsequently, we analyze the circumstances under which these conditions can be met. We show that under a suitable probability model of the calibration parameters, the sufficient conditions for recovery are satisfied with a higher probability with increasing number of coils. This means that, in practice, parallel imaging is likely to perform better than single imaging recovery.
@@ 4.2 A sufficient condition
(Uniqueness Results) Let $\mathbf{X}_{0}\in\mathbb{C}^{kN\times C}$ and $S\subset[N]$ be arbitrary. Consider a linear map A from $\mathbb{C}^{kN}$ to $\mathbb{C}^{L}$ and noisy measurements $\textbf{Y}=\textbf{A}\mathbf{X}_{0}+\boldsymbol{\mathcal{N}}$ with $\|\boldsymbol{\mathcal{N}}\|_{F}\leq\sigma$.
@@
Moreover, suppose that there exists $\mathbf{V}\in\mathbb{C}^{L\times C}$ such that the approximate dual certificate $\boldsymbol{\mathcal{Y}}=\textbf{A}^{*}\textbf{V}$ satisfies the following conditions
[…]
If $\rho=\theta+\frac{\eta\beta}{1-\delta}<1$, $\mu=\frac{\sqrt{1+\delta}}{1-\delta}$, then any solution $\textbf{X}^{*}$ of the problem (16) obeys
@@
The derivation of the lemma is a direct application of Lemma 7 of [20] which relies on the sub-differential of $\partial(\|\cdot\|_{(1,2)})$. Given $\textbf{Z}\in\mathbb{C}^{kN\times C}$, the sub-differential $\partial_{\textbf{Z}}(\|\cdot\|_{(1,2)})$ being
@@ 4.3 Conditioning of $\mathbf{A}_{S}$
Next, we give the conditions on when our matrix $\mathbf{A}$ satisfies the condition $\|\textbf{A}_{S}^{*}\textbf{A}_{S}-\textbf{I}_{S}\|\leq\delta$. Recall our forward matrix $\mathbf{A}_{S}=\mathbf{\mathbf{F}}_{\Omega}\mathbf{\Phi}_{S}$, the challenges come from the fact that its rows are dependent. Instead of bounding the norm directly, we apply the triangular inequality and work on bounding the two terms in $\underbrace{\|\mathbf{\Phi}_{S}^{*}\mathbf{\textbf{F}}_{\Omega}^{*}\mathbf{\textbf{F}}_{\Omega}\mathbf{\Phi}_{S}-\mathbf{\Phi}_{S}^{*}\mathbf{\Phi}_{S}\|}_{(1)}+\underbrace{\|\mathbf{\Phi}_{S}^{*}\mathbf{\Phi}_{S}-\mathbf{I}_{S}\|}_{(2)}$ separately. Indeed, even when the full Fourier space is sampled, i.e., $\mathbf{F}_{\Omega}=\mathbf{F}$, the first term is zero and we would at the very least, require the condition (2) to be satisfied.
@@
For operator $\mathbf{\Phi}$ defined in (13) which follows the randomness assumptions 4.1,
@@
The probability is over the randomness in B. The proof of the Lemma is derived from Lemma 4.3 of [36], where the authors consider the case B to be fixed but the rows of $\mathbf{\Psi}_{S}$ are chosen uniformly at random with replacement from the DFT matrix.
Remark: Due to the symmetry of $\mathbf{\Psi}$ and B in constructing the forward matrix $\mathbf{\Phi}$ (13), the result of Lemma (4) also holds true when B is fixed but the rows of $\mathbf{\Psi}$ chosen uniformly from all possible row permutations of a fixed matrix.
@@
Note that both the bounds are useful only when B and $\mathbf{\Psi}$ are “flat”, i.e., $\mu^{\textbf{B}}_{max}\sim O(1)$ and $\mu^{\mathbf{\Psi}}_{max}\sim O(1)$.
The next lemma addresses a bound for the parameter $\beta$ in $\max\limits_{j\in|S^{c}|}\|\textbf{A}_{S}^{*}\textbf{A}_{S^{c}_{j}}\|_{2\rightarrow 2}\leq\beta$ in 3.
@@
Consider the forward matrix A defined in (13) and asssume that $\|\textbf{A}_{S}^{*}\textbf{A}_{S}-\textbf{I}_{S}\|_{2\rightarrow 2}\leq\eta$,

### v2

@@ 4.1 Assumptions and main theorem
Before presenting the main theorem, we first summarize the assumptions on the forward model in (25) and on the signals for easy reference.
  1. (A1).
    The rows of B are chosen uniformly at random without replacement from the rows of a matrix $\textbf{B}_{0}$ with $\textbf{B}_{0}^{*}\textbf{B}_{0}=\textbf{I}_{k}$. (Equivalently, the columns of $\textbf{B}_{0}$ form a tight frame.)
  2. (A2).
    $\mathbf{\Psi}$ is a known and fixed orthonormal basis with $\mathbf{\Psi}^{*}\mathbf{\Psi}=\mathbf{\Psi}\mathbf{\Psi}^{*}=\textbf{I}_{N}$.
  3. (A3).
    The rows of $\textbf{F}_{\Omega}$ are chosen uniformly at random without replacement from the DFT matrix F with $\textbf{F}^{*}\textbf{F}=\frac{N}{L}\textbf{I}_{N}$.
  4. (A4).
    We analyze two probability models/distributions for the sensitivity parameters:
@@
We briefly introduced some common choices of B and $\boldsymbol{\Psi}$ in Section 2, including the smoothness promoting polynomial or sinosudal basis for B and DCT or wavelet transformation for $\boldsymbol{\Psi}$. Here, we further comment on our choice of assumptions (A1)-(A4). (A1) and (A3) are regarding the sensing matrix. (A3) is a standard assumption in compressive sensing theory concerning the random sampling scheme for k-space/Fourier space measurements. For (A1), B is chosen uniformly at random from all possible row permutations of a fixed matrix $\textbf{B}_{0}$. The orthogonality of the columns of B is not affected by row permutations since $\textbf{B}_{i}^{*}\textbf{B}_{j}=\sigma(\textbf{B}_{i})^{*}\sigma(\textbf{B}_{j})$. With B being a matrix with randomly permuted rows, for any fixed h, the coil sensitivity $\mathbf{s}=\textbf{B}\textbf{h}$ will be randomly permuted copies of some underlying $\textbf{s}_{0}=\textbf{B}_{0}\textbf{h}$. On the other hand, if B were deterministic, any zero entries in $\textbf{s}=\textbf{B}\textbf{h}$ will make it impossible to recover uniquely the values of the image z at those locations. The randomness in B helps in providing an average case analysis for each coil. Assumption (A4) on H is critical for developing an average case analysis across coils to explain the improved results when more channels are included.
Furthermore, we consider the underlying image $\textbf{z}=\mathbf{\Psi}\textbf{x}$ to be fixed and only impose some probability distribution on $\textbf{h}_{i}$. That is to assume we are imaging a fixed object with varying coil sensitivities where the spanning coefficient of the coils follows a certain probability distribution. For the two choices on the probability models on H, they all ensure incoherence between the coils and that the coil sensitivities have the same energy. Observe first that, the incoherence between coils $\textbf{s}_{i}^{*}\textbf{s}_{j}=(\textbf{B}\textbf{h}_{i})^{*}(\textbf{B}\textbf{h}_{j})=\textbf{h}_{i}^{*}\textbf{B}^{*}\textbf{B}\textbf{h}_{j}=\textbf{h}_{i}^{*}\textbf{h}_{j}$ is small for $k$ large, and the coils all have the same energy $\textbf{s}_{i}^{*}\textbf{s}_{i}=(\textbf{B}\textbf{h}_{i})^{*}(\textbf{B}\textbf{h}_{i})=\textbf{h}_{i}^{*}\textbf{B}^{*}\textbf{B}\textbf{h}_{i}=\textbf{h}_{i}^{*}\textbf{h}_{i}=1$.
@@
Before moving on to stating the theorem, let us discuss the reason for imposing such a probability model on H. Consider the worst case scenario where all the $\textbf{h}_{i}$ are identical, then the measurements $\textbf{Y}=\textbf{A}\textbf{X}$ consist of identical columns. By including multiple channels, we are not gaining more information on X regardless of the number of coil used. The following proposition gives a formal statement of the worst case analysis. The proof follows exactly from Proposition 4.1 in [Eldar_Rauhut_2010].
@@
However, one would expect that including multiple channels will provide more information to aid the recovery of the shared underlying object of interest. In a pMRI setting, it has been shown that parallel imaging (pMRI) outperforms single-coil MRI in many cases, and improved incoherence in the sensing matrix is often used as an explanation for such improvement. To understand the role of including multiple channels, and from the worst case analysis, it suggests a need to impose some probability model on the sensing parameters in order to prove some average case results.
Next, we state our main theorem that provides a sufficient condition on robust and stable recovery.
Theorem 1
For any arbitrary subset $S\in[N]$ with $\text{card}(S)=n$ and $\textbf{X}_{0}:=\textbf{z}_{0}\otimes\textbf{H},\textbf{z}_{0}\in\mathbb{C}^{N\times 1},\textbf{H}\in\mathbb{C}^{k\times C}$, consider the linear map $\textbf{A}\in\mathbb{C}^{L\times kN}$ as defined in (25) that satisfies the randomness assumptions (A1)-(A4) and the noisy measurement $\textbf{Y}=\textbf{A}\textbf{X}_{0}+\boldsymbol{\mathcal{N}}$ with $\|\boldsymbol{\mathcal{N}}\|_{F}\leq\sqrt{C}\sigma$. The solution X to the $\ell_{1,2}$ minimization problem (28) satisfies
$\|\textbf{X}-\textbf{X}_{0}\|_{F}\leq C_{1}\|\mathcal{P}_{S^{c}}\textbf{X}_{0}\|_{1,2}+(C_{2}+C_{3}\sqrt{s})\sqrt{C}\sigma$ (32)
with probability $1-4N^{-\alpha}$ if
  1. (C1.)
    the number of measurement satisfies
    $L\geq C_{\alpha}kn{\mu^{B}_{max}}^{2}{\mu^{\Psi}_{max}}^{2}\log(knN),$
  2. (C2.)
    $\textbf{h}_{l}$ satisfies the assumption (A4) and the number of coils satisfies
    $C\geq C_{\alpha}\log(N)k.$
Here, $\mu^{B}_{max}:=\max\limits_{i,j}\sqrt{L}|\mathbf{B}_{i,j}|$, $\mu^{\Psi}_{max}:=\max\limits_{i,j}\sqrt{N}|{\boldsymbol{\Psi}}_{i,j}|$, which measure the incoherence of the rows of B and of $\boldsymbol{\Psi}$, respectively. The constant $C_{\alpha}$ grows linearly with respect to $\alpha$.
A detailed bound for the constants in the theorem can be found in the proofs in Section 6. Next, we use Lemma 1 to get rid of the $\sqrt{C}$ scaling in the error bound in (32).
Remark: The result provides a sufficient condition for stable and robust recovery in the presence of noise and sparse deficiency in the signal. It is important to emphasize that if the signal is truly sparse and there is no noise present, perfect reconstruction is achieved (i.e., the error is zero). Hence, increasing the number of the coils does not provide any additional benefit when the number of measurements is sufficiently large. Furthermore, to understand the strength of the statement, note that if $S$ is the true support of an n-sparse $\textbf{z}_{0}$, $\|\mathcal{P}_{S^{c}}\textbf{X}_{0}\|_{1,2}=0$, then the number of required measurements $L$ is of the order of the sparsity level of $\textbf{X}_{0}$, up to log factors (i.e., $kn={{\|P_{S}(\textbf{X}_{0})\|}_{0}}$). For the constants in the bound of (C1), it will be meaningful when the magnitudes of the entries in $\mathbf{B}$ and $\boldsymbol{\Psi}$ do not vary too much. Optimally, when the columns of $\mathbf{B}$ and $\boldsymbol{\Psi}$ are chosen from the DFT matrices, we optimally have $\mu^{B}_{max}=1$, $\mu^{\Psi}_{max}=1$.
Ideally, we would like to derive a proof for the setup when both B and $\mathbf{\Psi}$ are fixed, and the randomness comes from the support set $S$ of z alone. The result would be analogous to the conditioning of the random sub-matrices of a fixed sampling matrix, e.g. see Chapter 14 of [foucart_rauhut_2015]. Yet, we encountered some (at least to us) unsolvable difficulties in finding a realistic bound for the incoherence measure between the sub-blocks of the sampling matrix A. Nevertheless, we still provide a formal statement of the bound and some discussions in the Appendix 7.1.
@@
The proof of the main theorem follows a well-established proof outline of compressive sensing [foucart_rauhut_2015]. First, we refer to a sufficient condition from [flinth_2017] to ensure a robust and stable recovery. Subsequently, we analyze the circumstances under which these conditions can be met. We show in Lemma 8 that under assumptions (A1)-(A4), the sufficient conditions for recovery are satisfied with a higher probability as the number of coils increases. This means that, in practice, parallel imaging is likely to perform better than single imaging recovery.
@@ 4.2 A sufficient condition
We state below a sufficient condition for solving the general the $\ell_{1,2}$ minimization problem for recovering block sparse signals.
@@
(Uniqueness Results) Consider any matrix $\mathbf{X}_{0}\in\mathbb{C}^{kN\times C}$ and arbitrary set $S\subset[N]$. Let A be a linear operator from $\mathbb{C}^{kN}$ to $\mathbb{C}^{L}$, and the noisy measurements $\textbf{Y}=\textbf{A}\mathbf{X}_{0}+\boldsymbol{\mathcal{N}}$ with $\|\boldsymbol{\mathcal{N}}\|_{F}\leq\sigma$.
@@
Moreover, suppose that there exists a matrix $\mathbf{V}\in\mathbb{C}^{L\times C}$ such that the approximate dual certificate $\boldsymbol{\mathcal{Y}}:=\textbf{A}^{*}\textbf{V}$ satisfies the following conditions
$\|\mathcal{P}_{S}\boldsymbol{\mathcal{Y}}-\text{sgn}(\textbf{X}_{0})\|_{F}\leq\eta,\max\limits_{j\in|S^{c}|}\|\mathcal{P}_{S^{c}_{j}}\boldsymbol{\mathcal{Y}}\|_{F}\leq\theta\text{ and }\|\textbf{V}\|_{F}\leq\tau\sqrt{s}.$ (34)
Define $\rho=\theta+\frac{\eta\beta}{1-\delta}$, $\mu=\frac{\sqrt{1+\delta}}{1-\delta}$. If $\rho<1$, then the solution $\textbf{X}^{*}$ to the problem (27) obeys
@@
The derivation of the lemma is a direct application of Lemma 7 of [flinth_2017] which relies on the sub-differential of $\partial(\|\cdot\|_{(1,2)})$. Given $\textbf{Z}\in\mathbb{C}^{kN\times C}$, the sub-differential $\partial_{\textbf{Z}}(\|\cdot\|_{(1,2)})$ is the set
@@
In the following sections, we will prove that with the assumptions we have made, the parameters $\delta$ and $\beta$ will be small with high probability. We will also construct a dual certificate $\boldsymbol{\mathcal{Y}}$ with small $\eta$, $\theta$ and $\tau$. Most of the proof procedure follows well-established methods in the compressive sensing literature, such as those described in [ling_strohmer_2015] and [flinth_2017].
@@ 4.3 Conditioning of $\mathbf{A}_{S}$
Next, we give the conditions on when our measurement matrix $\mathbf{A}$ satisfies the condition $\|\textbf{A}_{S}^{*}\textbf{A}_{S}-\textbf{I}_{S}\|\leq\delta$. Recall our forward matrix $\mathbf{A}_{S}=\mathbf{\mathbf{F}}_{\Omega}\mathbf{\Phi}_{S}$, the challenges come from the fact that its rows are dependent. Instead of bounding the norm directly, we apply the triangular inequality and work on bounding the two terms in $\underbrace{\|\mathbf{\Phi}_{S}^{*}\mathbf{\textbf{F}}_{\Omega}^{*}\mathbf{\textbf{F}}_{\Omega}\mathbf{\Phi}_{S}-\mathbf{\Phi}_{S}^{*}\mathbf{\Phi}_{S}\|}_{(1)}+\underbrace{\|\mathbf{\Phi}_{S}^{*}\mathbf{\Phi}_{S}-\mathbf{I}_{S}\|}_{(2)}$ separately. Indeed, even when the full Fourier space is sampled, i.e., $\mathbf{F}_{\Omega}=\mathbf{F}$, the first term (1) is zero and we would at the very least, require the condition (2) to be satisfied.
@@
For operator $\mathbf{\Phi}$ defined in (23) which follows the assumptions induced by $\boldsymbol{\Psi}$ and B from (A1)-(A4) 4.1,
@@
The probability is over the randomness in B. The proof of the Lemma is derived from Lemma 4.3 of [ling_strohmer_2015], where the authors consider the case B to be fixed but the rows of $\mathbf{\Psi}_{S}$ are chosen uniformly at random with replacement from the DFT matrix.
Remark: Due to the symmetry of $\mathbf{\Psi}$ and B in constructing the forward matrix $\mathbf{\Phi}$ (23), the result of Lemma 4 also holds true when B is fixed but the rows of $\mathbf{\Psi}$ chosen uniformly from all possible row permutations of a fixed matrix.
@@
Combining Lemma 4 and Lemma 5, we have a probabilistic bound for $\|\textbf{A}_{S}^{*}\textbf{A}_{S}-\textbf{I}_{S}\|_{2\rightarrow 2}$. Note that both the bounds are useful only when B and $\mathbf{\Psi}$ are “flat”, i.e., $\mu^{\textbf{B}}_{max}\sim O(1)$ and $\mu^{\mathbf{\Psi}}_{max}\sim O(1)$.
The next lemma provides a bound for the parameter $\beta$ in $\max\limits_{j\in|S^{c}|}\|\textbf{A}_{S}^{*}\textbf{A}_{S^{c}_{j}}\|_{2\rightarrow 2}\leq\beta$ in Lemma 3.
@@
Consider the forward matrix A defined in (23) and assume that $\|\textbf{A}_{S}^{*}\textbf{A}_{S}-\textbf{I}_{S}\|_{2\rightarrow 2}\leq\eta$,